Calculate the average molecular speed of O_2 at 300 K in m/s.
1
Expert's answer
2020-06-22T13:34:20-0400

vrms = (3RT / m0)1/2 = (3 × 8.314 × 300 / 0.032 )1/2 = 483 (m / s)